OpenStudy (anonymous):

A circle has its center at (1, 4) and a radius of 2 units. What is the equation of the circle? (1 point) Select one: a. (x + 2)2 + (y + 4)2 = 2 b. (x - 1)2 + (y - 4)2 = 4 c. (x + 1)2 + (y - 4)2 = 4 d. (x - 1)2 + (y - 4)2 = 2

OpenStudy (campbell_st):

the general form is \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre and r is the radius substitute your information to get the answer

OpenStudy (campbell_st):

the choices have evaluated the radius squared.

OpenStudy (anonymous):

so it would be D right?

OpenStudy (anonymous):

why d ?

OpenStudy (campbell_st):

remember is radius squared 2^2 =

OpenStudy (anonymous):

oh so b

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

ok thank you

OpenStudy (campbell_st):

thats it